OpenStudy (anonymous):

Jerry skipped some steps and made and error in his solution. Explain the error in Jerry's solution below. Give the correct solution(s). x\left( {x + 6} \right) = - 5 x = - 5 \begin{array}{l}x + 6 = - 5\\x = - 11\end{array}

OpenStudy (misty1212):

\[x\left( {x + 6} \right) = - 5 \]

OpenStudy (misty1212):

just trying to figure out what you wrote

OpenStudy (misty1212):

\[x(x+6)=-5\\ x=-5\\ x+6=-5, x=-11\] something like that?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

the error is that if \(x(x+6)=-5\) then it does not follow that one of those factors is \(-5\) we can only use that if it is set equal to zero i.e. if the product of two factors are zero, one of the factors must be zero